OpenStudy (anonymous):

Which is an equation of a circle with center (-5, -7) that passes through the point (0, 0)?

OpenStudy (jdoe0001):

 what's the radius of it anyway? \(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -5}}\quad ,&{\color{blue}{ -7}})\quad &({\color{red}{ 0}}\quad ,&{\color{blue}{ 0}}) \end{array}\qquad d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2} \\ \quad \\\\ \quad \\ (x-{\color{brown}{ h}})^2+(y-{\color{blue}{ k}})^2=r^2\qquad center\ ({\color{brown}{ h}},{\color{blue}{ k}})\quad radius=r\)